OpenStudy (shelld93):

A student bought 1 Blu-Ray and 2 DVD movies for a total of $90, while a second student spent $171 to purchase 5 DVDs and 1 Blu-Ray. If B is the cost of a Blu-Ray in dollars and D is the cost of a DVD in dollars: a) Write two (2) equations that relate the number of Blu-Rays and DVDs to their cost. b)Solve the above equations simultaneously to find the cost of a Blu-Ray and the cost of a DVD. c) Check your solution.

OpenStudy (johnweldon1993):

Let Blue Ray = b and let DVD = d If they bought 1 blue ray and 2 dvds for 90 that looks like b + 2d = 90 Now if another student bought 5 dvds and 1 blue ray for 171....that is b + 5d = 171 So we have our 2 equations \[\large b + 2d = 90\]\[\large b + 5d = 171\]

OpenStudy (shelld93):

So far I have only dealt with one letter in a questions, how do I go about solving these ?

OpenStudy (johnweldon1993):

Oh okay...well then this is good that we have 2 equations then.... When you have more than 1 variable...(2 here)you need that many equations (2) to solve for it... Alright....so choose 1....elimination method (easier here) or substitution method

OpenStudy (shelld93):

uhm i dont know the difference so do the easier elimination

OpenStudy (johnweldon1993):

Lol alright...and both are ways to solve systems of equations here...(I'll do substitution later so you can see it too :) So elimination...the process of eliminating 1 variable so you can solve for the other... Start with both equations \[\large b + 2d = 90\]\[\large b + 5d = 171\]

OpenStudy (johnweldon1993):

Now subtract them... Why? because if you look...if you were to subtract the 2 equations up there....the 'b' variable would go away (eliminated) because b - b is 0

OpenStudy (shelld93):

what are we subtracting ?

OpenStudy (johnweldon1993):